OpenStudy (rock_mit182):

Evaluate following two compositions: a: (f ° g) (x) b: (h° g) (x) f(x) = 2x+5; g(x) = x^2-3; h(x) = (7-x)/(3)

OpenStudy (rock_mit182):

f ((7-x)/(3) ) by definition (f°h)(x) = ((7-x)/(3) )^2 -3

OpenStudy (rock_mit182):

\[(f ° h)(x) = (\frac{ 7-x }{ 3 } )^{2}-3\]

OpenStudy (rock_mit182):

h (g(x)) by definition h (x^2-3)

OpenStudy (rock_mit182):

\[(h ° g)(x) = \frac{ 7-x ^{2}-3 }{ 3} \]

OpenStudy (rock_mit182):

(-x^2+4)/(3)

OpenStudy (rock_mit182):

@iPwnBunnies

OpenStudy (aum):

a: (f ° g) (x) = f( g(x) ) f(x) = 2x+5; g(x) = x^2-3. f( g(x) ) = 2(x^2-3) + 5 = 2x^2 - 6 + 5 = 2x^2 - 1

OpenStudy (rock_mit182):

A = (f ° g)(x) = f(g(x)) by definition \[f(x ^{2}-3)\] \[2(x ^{2}-3) +5\]

OpenStudy (rock_mit182):

\[(f ° g)(x) = 2x ^ {2} -6 +5 -->2x ^{2}-1\]

OpenStudy (aum):

correct. Do the next one the same way. b: (h° g) (x)

OpenStudy (aum):

b: (h° g) (x) = h( g(x) ) g(x) = x^2-3; h(x) = (7-x)/(3) (h° g) (x) = h( g(x) ) = h( x^2-3 ) = ( 7 - (x^2-3) ) / 3 = (7 - x^2 + 3) / 3 = (10-x^2)/3

OpenStudy (rock_mit182):

b: (h°g)(x) = h(g(x)) by definition h(x^2-3) \[\frac{ 7-(x ^{2}-3) }{ 3 } = \frac{ 10-x ^{2} }{ 3 }\]

OpenStudy (aum):

correct.

OpenStudy (rock_mit182):

THanks a lot, you have been helpfull

OpenStudy (aum):

you are welcome.

OpenStudy (rock_mit182):

\[h(x ^{2}-3)\]

OpenStudy (rock_mit182):

\[f ^{-1}(x) = 2y +5\]

OpenStudy (rock_mit182):

x= 2y + 5 -5 +x = 2y

OpenStudy (rock_mit182):

\[-\frac{ 5 }{ 2 }+\frac{ x }{ 2 } = y\]

OpenStudy (rock_mit182):

\[f ^{-1}(x) = \frac{ x }{ 2 }- \frac{ 5 }{ 2 }\]

OpenStudy (rock_mit182):

\[h ^{-1}(x) =\]

OpenStudy (rock_mit182):

\[x = \frac{ 7-y }{ 3 }\]

OpenStudy (rock_mit182):

\[3x=7-y\]

OpenStudy (rock_mit182):

\[3x-7=-y\]

OpenStudy (rock_mit182):

\[-3x+7 = y\]

OpenStudy (rock_mit182):

\[h ^{-1}(x) = -3x+7\]

OpenStudy (rock_mit182):

\[(2x+5)- (\frac{ 7-x }{ 3 }) = 2x+5 -\frac{ 7 }{ 3 }+\frac{ x }{3 }\]